find 2 numbers thats sum is 28 and a product of 147
Found 2 solutions by richwmiller, drk:
Answer by richwmiller(17219)

a+b=28
ab=147
a=28-b
(28-b)*b=147
28b-b^2=147
0=b^2-28b+147
(b-7)(b-21)=0
b=21 or b=7
a=7 or 21

Answer by drk(1908)

Let X and Y be the 2 numbers.
(i) X + Y = 28
(ii) XY = 147
--
Step 1 - solve (i) for X and we get
(iii) X = 28 - Y
step 2 - substitute (iii) into (ii) to get
(iv) (28-Y)(Y) = 147
step 3 - distribute to get
(v) -y^2 + 28y = 147
step 4 - set = 0 and factor as
(vi) -y^2 + 28y -147 = 0
factoring, we get
(vii) -(y-21)(y-7) = 0
solving each for y, we get
y = 21, y = 7
So, we get two answers:
x = 7, y = 21
x = 21, y = 7.